the coastline of california is 46 miles longer than twice the length of louisianas coastline it is also 443 miles longer than louisianas coastline find the lengths of the coastlines of california and louisiana
C= CAL COAT
L= LOUISIANA COAST
C = 2L+46
C=L+443
2L+46 = L+443
L=443-46
L = 397 MILES LOUISIANA COAST

C = 2(397)+46
C =840 MILES CAL. COAST

Answer by chiexpert(48)


C= CAL COAT
L= LOUISIANA COAST
C = 2L+46
C=L+443
2L+46 = L+443
L = 397 MILES LOUISIANA COAST

C = 2(397)+46
C =840 MILES CAL. COAST